# NAAQA: A Neural Architecture for Acoustic Question Answering

### v1

@@ Appendix A Network Optimization
512 94.2 ±0.23 91.5 ±0.54 91.4 ±0.39 $\sim$1.94 M
1,024 93.5 ±0.93 91.3 ±0.53 91.2 ±0.46 $\sim$2.46 M
256 92.8 ±0.70 91.3 ±0.49 91.1 ±0.07 $\sim$1.68 M
4,096* 94.1 ±0.62 91.2 ±0.33 90.9 ±0.21 $\sim$5.61 M
2,048 92.7 ±1.03 87.4 ±5.29 86.5 ±5.69 $\sim$3.51 M
TABLE VI: Impact of the number of GRU text-processing units $G$. The interleaved extractor is used with hyper-parameters from the initial configuration described in Section 5.3 except for $G$. The row marked with * correspond to the initial value for $G$. The number of GRU unit have a strong impact on the number of parameters. Setting $G=512$ actually increase accuracy while reducing the number of parameters by a factor of $\sim$2.8.
Nb (C) Hidden Overall Accuracy (%) Nb of
Filter Layer (H) Train Val Test Parameters
256 1024 93.8 ±0.49 91.4 ±0.36 91.6 ±0.08 $\sim$1.64 M
512* 1024* 94.2 ±0.23 91.5 ±0.54 91.4 ±0.39 $\sim$1.94 M
128 1024 94.3 ±0.48 91.3 ±0.45 91.2 ±0.52 $\sim$1.50 M
128 256 93.2 ±0.61 91.0 ±0.25 91.1 ±0.55 $\sim$1.35 M
256 256 93.1 ±0.56 91.0 ±0.66 91.0 ±0.60 $\sim$1.40 M
128 512 92.9 ±0.41 90.9 ±0.34 90.8 ±0.29 $\sim$1.40 M
256 512 93.4 ±0.67 90.9 ±0.55 90.7 ±0.28 $\sim$1.48 M
512 512 77.4 ±21.55 76.2 ±21.30 76.1 ±21.25 $\sim$1.65 M
TABLE VII: Impact of the number of filters $C$ and hidden units $H$ in the classifier. The interleaved extractor is used with hyper-parameters from the initial configuration described in Section 5.3 except for $C$, $H$ and $G=512$. The row marked with * correspond to the initial value for $C$ and $H$. Classifier parameters have a small impact on both accuracy and the number of parameters. Setting $C=256$ gives similar accuracy and reduce the number of parameters by 0.3M compared to the initial parameters. The configuration $C=512$ and $M=512$ get stuck in local minimums often which explain the large STD.
@@
4 128 94.0 ±0.54 91.3 ±0.40 91.7 ±0.07 $\sim$1.64 M
4 64 93.7 ±0.04 91.0 ±0.26 91.1 ±0.04 $\sim$0.83 M
3 128 93.5 ±0.35 90.6 ±0.81 90.6 ±0.68 $\sim$1.35 M
3 64 92.8 ±0.90 90.5 ±0.61 90.3 ±0.61 $\sim$0.73 M
2 128 93.4 ±0.15 90.7 ±0.11 90.2 ±0.14 $\sim$1.05 M
2 64 92.0 ±0.39 89.6 ±0.25 89.6 ±0.14 $\sim$0.62 M
3 32 91.6 ±0.49 88.8 ±0.25 88.6 ±0.24 $\sim$0.51 M
4 32 91.8 ±0.58 89.3 ±1.05 88.6 ±1.36 $\sim$0.55 M
2 32 90.2 ±1.05 86.5 ±0.77 86.7 ±0.65 $\sim$0.47 M
1 128 92.0 ±0.22 87.1 ±0.35 86.3 ±0.13 $\sim$0.76 M
1 64 91.1 ±1.55 86.0 ±1.13 86.0 ±1.31 $\sim$0.51 M
1 32 87.5 ±0.25 81.1 ±0.46 81.4 ±0.26 $\sim$0.42 M
TABLE VIII: Impact of reducing the number of Resblock $J$ and the number of filters $M$ in each Resblocks. The interleaved extractor is used with hyper-parameters from the initial configuration described in Section 5.3 except for $J$, $M$, $C=256$, $H=1024$ and $G=512$. With $J=4$ and $M=64$, the network reaches great accuracy with only 0.67M parameters (vs 5.61M for the initial configuration). The accuracy drop when $J<3$ and $M=32$. The configuration $J=3$ and $M=128$ get stuck in local minimums often which explain the large STD.
Nb (K) Projection (P) Accuracy (% ± 0.60) Nb of
Block Size Train val Test Parameters
3 32 93.7 ±0.34 91.5 ±0.03 91.5 ±0.15 $\sim$0.81 M
3 64 93.7 ±0.04 91.0 ±0.26 91.1 ±0.04 $\sim$0.83 M
3 128 93.1 ±0.24 91.0 ±0.66 91.1 ±0.36 $\sim$0.87 M
3 93.5 ±0.30 91.1 ±0.24 90.8 ±0.29 $\sim$0.81 M
2 92.9 ±0.39 89.9 ±0.13 90.0 ±0.13 $\sim$0.80 M
2 32 92.9 ±0.72 89.7 ±0.62 89.9 ±0.55 $\sim$0.81 M
2 128 91.9 ±1.80 89.5 ±0.75 89.5 ±1.03 $\sim$0.87 M
2 64 91.9 ±1.14 89.4 ±0.57 89.4 ±0.58 $\sim$0.83 M
4 32 93.4 ±0.47 86.9 ±2.43 83.1 ±0.07 $\sim$0.83 M
4 64 89.0 ±0.00 81.6 ±0.00 80.8 ±0.00 $\sim$0.86 M
4 91.3 ±2.00 83.5 ±6.44 79.9 ±4.09 $\sim$0.85 M
4 128 88.2 ±2.37 77.3 ±2.70 74.8 ±2.62 $\sim$0.90 M
TABLE IX: Impact of the number of blocks $K$ in the parallel feature extractor and its projection size $P$. The interleaved extractor is used with the hyper-parameters $J=4$, $M=64$, $C=256$, $H=1024$, $G=512$. More blocks in the extractor results in more downsampling of the feature maps. Too much downsampling ($K=4$) negatively affects accuracy.

### v2

@@ Acknowledgment
512 95.5 ±0.28 90.1 ±0.37 79.6 ±0.30 1.94 M
256 94.3 ±0.51 90.5 ±0.30 79.5 ±0.05 1.68 M
1,024 94.2 ±0.31 89.7 ±0.18 79.0 ±0.14 2.46 M
4,096 94.8 ±0.25 89.9 ±0.25 78.9 ±0.23 5.61 M
2,048 95.8 ±0.14 90.0 ±0.05 78.7 ±0.04 3.51 M
TABLE VIII: Impact of the number of GRU text-processing units $G$. The interleaved extractor is used with hyper-parameters from the initial configuration described in Section 5.3 except for $G$. The number of GRU unit has a strong impact on the number of parameters. Setting $G=512$ actually increases accuracy while reducing the number of parameters by a factor of $\sim$2.8. The rows are ordered by test accuracy.
@@
4 128 94.9 ±0.91 89.9 ±0.39 79.1 ±0.56 1.94 M
3 128 95.0 ±0.67 89.4 ±0.42 78.4 ±0.25 1.64 M
2 128 93.6 ±0.40 86.9 ±0.61 77.3 ±0.88 1.35 M
4 64 92.8 ±0.95 87.6 ±0.98 77.0 ±0.66 1.11 M
3 64 91.8 ±1.36 86.2 ±0.77 76.4 ±0.65 1.01 M
2 64 90.8 ±0.75 84.5 ±0.64 75.0 ±0.63 0.90 M
4 32 88.1 ±1.35 83.1 ±1.08 73.6 ±1.18 0.82 M
3 32 86.3 ±1.39 80.6 ±1.54 71.5 ±1.54 0.78 M
2 32 84.4 ±1.36 76.8 ±2.02 68.4 ±1.59 0.74 M
1 128 89.0 ±1.80 75.3 ±2.65 67.2 ±2.43 1.05 M
1 64 84.3 ±4.61 69.6 ±4.61 62.1 ±3.41 0.79 M
1 32 67.0 ±3.58 57.3 ±2.07 52.9 ±1.50 0.69 M
TABLE IX: Impact of reducing the number of Resblock $J$ and the number of filters $M$ in each Resblocks. The interleaved extractor is used with hyper-parameters from the initial configuration described in Section 5.3 except for $J$, $M$, $C=256$, $H=1024$ and $G=512$. With $J=4$ and $M=64$, the network reaches great accuracy with only 0.67M parameters (vs 5.61M for the initial configuration). The accuracy drops when $J<3$ and $M=32$. The configuration $J=3$ and $M=128$ get stuck in local minimums often which explain the large STD. The rows are ordered by test accuracy
Tables VIII and IX report the intermediate results of the NAAQA network complexity study on CLEAR2.